### v2

Abstract
Outlier detection is an important data mining tool that becomes particularly challenging when dealing with nominal data. First and foremost, flagging observations as outlying requires a well-defined notion of nominal outlyingness. This paper presents a definition of nominal outlyingness and introduces a general framework for quantifying outlyingness of nominal data. The proposed framework makes use of ideas from the association rule mining literature and can be used for calculating scores that indicate how outlying a nominal observation is. Methods for determining the involved hyperparameter values are presented and the concepts of variable contributions and outlyingness depth are introduced, in an attempt to enhance interpretability of the results. The proposed framework is evaluated on both synthetic and real-world data sets, demonstrating comparable performance to state-of-the-art frequent pattern mining algorithms and even outperforming them in certain cases. The ideas presented can serve as a tool for assessing the degree to which an observation differs from the rest of the data, under the assumption of sequences of nominal levels having been generated from a Multinomial distribution with varying event probabilities.
keywords:
Outlier detection, Nominal data, Association rule mining, Contingency tables, Multinomial distribution
pacs:
MSC Classification]62H30
1 Introduction
Outlier detection (also known as anomaly detection) is a crucial aspect of data analysis that has received increasing attention by data mining researchers. The wide range of uses of outlier detection in various domains, such as in cybersecurity [1, 18], sociology [39], healthcare [45] and finance [35], has driven the development of numerous algorithms designed to detect anomalies within a data set. Most of these algorithms are restricted to just one data type and that is continuous data. However, categorical data, that is binary, ordinal and nominal data [22], is prevalent in real-world data sets, yet very few outlier detection methods can handle this data type and authors do not always make the distinction between purely nominal and ordinal variables. Outliers in data sets with nominal variables may indicate a serious offence or a critical situation. For instance, unusual combinations of symptoms, diagnoses and demographic characteristics of patients could be indications of rare conditions in healthcare monitoring or uncommon combinations of claim types, policies and accident locations may hint fraud in insurance claims. This demands the development of outlier detection methods tailored for nominal data.
In this paper, we focus on nominal variables, including binary features. For completeness, we present some of the related literature below. Since no standard method exists, we provide a brief overview of the various approaches that have been proposed. For an extensive review of outlier detection algorithms for categorical data, in the sense of both nominal and ordinal data, we refer the reader to [44].
One approach to outlier identification assuming an underlying discrete distribution, is extending ideas from the continuous to the discrete domain. This would typically require robust estimation of the parameters of some assumed underlying distribution, computation of robust distances and finally comparing the distances to a cutoff value [see 43, 37, 38, for examples of robust estimators]. The cutoff value is typically the $1-\alpha$ quantile of the distribution of the distances, assuming validity of the hypothesised data generating mechanism. In the discrete distribution setting, this is generally a hard task that this is not possible unless some model is assumed to have generated the counts in a contingency table. Log-linear models are widely used for this purpose and one can then infer what the expected proportions should be in a robust way [see for example 27, 28, 11]. However, these methods can become impractical when the number of nominal variables is very large and the analysis of a high-dimensional contingency table is required. The sparsity of the resulting contingency tables increases together with the dimensionality, which is something that needs to be accounted for as well.
Another notable approach to outlier detection for nominal data is using proximity-based methods. These methods seek to detect outliers by generating a distance matrix between observations and finding the subjects that are furthest away from the rest of the points in the data set. Examples of such methods include [6] and [32]. However, distances in the categorical space often lack meaningful interpretation and implicitly make an equidistance assumption. As an example, the Hamming distance treats all mismatches equally. [42] introduced a distance metric for detecting outliers in categorical data; however, its application is restricted to $2\times 2$ contingency tables. Furthermore, these methods are very impractical in the presence of a large number of nominal variables, leading to a high computational cost. Information-theoretic methods have also been proposed for outlier identification of nominal data [30, 47, 48]. Such methods rely on some information-theoretic measure of data disorder, such as the entropy, and typically employ a local search algorithm to optimise an objective function. As a result, these methods rely heavily on heuristics and the optimisation process can be very costly. These problems are also common in recently proposed deep learning methods [for example 13, 14].
One different approach to anomaly detection for nominal data is the rule-based approach that makes use of ideas from the association rule mining literature [4]. To the best of our knowledge, the first algorithm that adopted this mindset for dealing with the problem of outlier detection in nominal data was the Link-based Outlier and Anomaly Detection in Evolving Data sets (LOADED) algorithm of [19]. In fact, LOADED was developed for identifying anomalies in mixed-type data (that is for a combination of nominal and continuous features). An improvement of this algorithm in terms of computational cost was then presented by [36], with further ameliorations being proposed by [24] for the nominal outlier identification part. More precisely, [24] introduced the Outlier Detection for Mixed Attribute Datasets (ODMAD) algorithm. The primary concern with the way ODMAD deals with nominal features is that it treats outliers in the nominal domain in a way that disregards the dimension of the subspace in which these are found. Furthermore, the implementation of ODMAD involves certain threshold parameters, the values of which need to be defined by the user, without providing guidelines of how these could be chosen in a data-driven manner. In a similar fashion, several frequent pattern mining methodologies have been employed for quantifying nominal outlyingness, such as the Frequent Pattern Outlier Factor (FPOF) [21] and the Frequent Pattern Isolation (FPI) [26] algorithms. However, these methods also suffer from a lack of intuition regarding the choice of their associated hyperparameter values and do not take into account that longer sequences are inherently less frequent than their subsets. Moreover, they completely overlook whether an event is rare by nature. For instance, consider a rare disease that affects only a small proportion of the population but is known and not particularly noteworthy. Frequent pattern mining algorithms will most likely flag subjects suffering from that disease as potential outliers solely due to their low frequency within a data set of clinical records, even though these are completely normal and expected cases. This underscores the need for a framework that can incorporate such contextual information and produce results that align with the underlying ground truth.
The rest of the paper is organised as follows: we define the concept of nominal outlyingness by adopting a rule-based mindset in Section 2 and we give a brief introduction to key concepts from the area of association rule mining. We then propose a score that quantifies abnormal behaviour in Section 3, focusing on the rationale behind its formulation and how the associated hyperparameters can be determined from the data by making the assumption of an underlying data generating mechanism. Some mathematical properties of our proposed score are included in Section 4. The efficacy of our method is illustrated via a series of simulations on synthetic and real-world data in Section 5, where our proposal is benchmarked against two state-of-the-art frequent pattern mining algorithms. We finally make some concluding remarks in Section 6.
2 Outlyingness in discrete space
In this Section we extend concepts of multivariate outlier detection for continuous data to discrete spaces. We define the concept of nominal outlyingness and give a brief introduction to basic notions from the association rule mining.
2.1 Definition of nominal outlyingness
One may find several definitions of outlyingness for continuous observations in the literature. All these definitions typically rely on some assumptions, for example related to the data generating mechanism [17, 7], or the location of each observation relative to the rest of the points in the feature space [e.g. 10, 33]. If nominal variables appear in a data set, using the location of points becomes difficult to work with and looking at distances in the categorical space is not ideal for reasons previously stated in Section 1. As we are dealing with point masses instead of densities, only a fixed number of values can occur, meaning that location may no longer be used as a criterion to assess anomalous behaviour. However, one can make use of the ideas related to the identification of outliers with respect to the mechanism that has been used to generate the data. In fact, all we have to assume is that the data has been generated from some underlying discrete distribution. Then, we turn our attention to observations of abnormally high or low mass with respect to this discrete distribution, just like we would suspect points of low density to be potential outliers in the continuous setting. This leads to two possible definitions of nominal outlyingness; one for observations whose nominal values are too frequent and one for too infrequent values, always with respect to an assumed discrete random variable. The notion of frequency is described in detail later in the section.
Notice that looking at both highly frequent and highly infrequent values is rather unsound; this is easy to realise by considering a toy example of one binary variable generated by a Bernoulli distribution for which the success probability has been significantly misspecified. In such a case, one of the two values will be treated as highly infrequent. However, this implicitly means that the other value will be highly frequent. Therefore, one needs to specify whether their interest lies in highly frequent or highly infrequent values, that is basically what they wish to define as a nominal outlier.
Extending these ideas to multiple nominal features requires taking combinations of their values into account. One potential downside of such an approach is that in the presence of multiple nominal variables taking a large number of values, it is very likely for almost every single combination of values to appear fewer times than expected. As a result, our definition of nominal outlyingness considers sequences of nominal values, weighing their contribution to outlyingness in a way that is inversely proportional to the sequence length. This is done in order to ensure that, if interest lies in highly infrequent sequences, combinations of values are seen with greater suspicion if they are shorter in length, as they are typically expected to be more abundant than their supersets in a data set. Based on these considerations, we can now define nominal outlyingness as follows:
Definition 1.
Nominal outlyingness is the degree to which a sequence of nominal values and its subsets are markedly different from an assumed discrete distribution that has generated them.
In order to introduce a framework for quantifying nominal outlyingness as given in Definition 1, we borrow some ideas from the association rule mining literature. We briefly outline some basic concepts in the following subsection.
2.2 Association rule mining background
Association rule mining is a method used to discover relationships of interest between sets of data items. The work of [4] was one of the first to focus on generating association rules between items in a large database of customer transactions and has paved the way for related concepts to be used in market basket analysis [46], customer segmentation [40] and other application areas. Below, we present some basic notions used in association rule mining. We later make use of these concepts in the development of a suitable score that quantifies how outlying a nominal observation is.
We start by defining an itemset as a set of items which occur together. Here, item refers to the value any variable takes for each observation. Since we are working with nominal data, an itemset is simply a sequence of categorical levels. We denote an itemset by $d$ and its length by $\lvert d\rvert$. Notice that $1\leq\lvert d\rvert\leq p$, where $p$ is the number of variables and $\lvert d\rvert=1$ refers to one level of a single nominal variable. Finally, we define the support of a an itemset $d$, denoted by $\operatorname{supp}(d)$, as the number of times the itemset $d$ appears in the data.
Given a large data set, some itemsets may appear more often than others. In an outlier identification problem, where the focus is on highly frequent or highly infrequent occurrences, determining whether an itemset is frequent or infrequent is of a major importance. An itemset can be considered as frequent when its support exceeds a minimum support threshold value, typically denoted by $\sigma$; otherwise it is infrequent. As we will see, this threshold value may differ according to the itemset that is being looked at, therefore we use the notation $\sigma_{d}$ instead, to denote that the threshold value is specific to an itemset $d$. Determining $\sigma_{d}$ is one of the most challenging problems in the association rule mining literature; this is something that we deal with later in the paper.
Finally, we introduce the Apriori principle and the related concept of support-based pruning. According to the Apriori principle [3], if an itemset is frequent, then all of its subsets must be frequent as well. Equivalently, if an itemset is infrequent, then all its supersets will also be infrequent. A direct implication of the Apriori principle is that one may not consider subsets or supersets of itemsets that are already found to be frequent or infrequent, respectively, something known as support-based pruning. The reason why support-based pruning is used is to avoid redundant computations and hence reduce the computational complexity of the search by trimming the space of itemsets that are being examined.
3 Proposed framework
The problem of outlier identification is generally complex in the sense that it is an unsupervised problem and thus, we can not be certain of whether an observation is truly outlying. However, what we can do is measure outlyingness with respect to a given scheme. By doing so, not only can we deduce how likely an observation is to be an outlier, but we can also compare observations in terms of their outlyingness. In this Section, we introduce a score that quantifies nominal outlyingness given Definition 1. We explain the rationale behind our score formulation and describe how values for the involved hyperparameters can be chosen. We finally introduce the concepts of nominal outlyingness depth and the contribution matrix which can be used for meta-analysis purposes.
3.1 Score of nominal outlyingness
Let us assume that we are working with a dataset $\mathcal{D}$ consisting of $p$ nominal variables $\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{p}$. We denote the number of levels for each variable $\boldsymbol{X}_{i}$ by $\ell_{i}$ and assume without loss of generality that its levels are encoded by the natural numbers $1,\ldots,\ell_{i}$ $(i=1,\ldots,p)$. Having defined some basic notions from the association rule mining literature, we treat each entry in a contingency table as an individual itemset. The total number of contingency tables is given by $12^{p}-$ 1 and the total number of itemsets that one can encounter is $\prod_{i=1}^{p}\left(\ell_{i}+1\right)-1$, assuming that an individual nominal feature $\boldsymbol{X}_{i}$ is a $(\ell_{i}\times 1)$-dimensional contingency table. We further define $\boldsymbol{\pi}_{i}$ to be the $\ell_{i}$-dimensional vector including the expected proportions for each level of $\boldsymbol{X}_{i}$ $(i=1,\ldots,p)$.
[…]
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r}},$ (1)
$r>0,\ i=1,\dots,n,$
if we are interested in highly infrequent itemsets. Recall that $\operatorname{supp}(d)$ refers to the support of itemset $d$, as defined in Section 2. Otherwise, for highly frequent itemsets, the score of nominal outlyingness becomes:
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin[0,\sigma_{d}),\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\operatorname{supp}(d)}{\sigma_{d}\times\left(\operatorname*{\operatorname{MAXLEN}}-\lvert d\rvert+1\right)^{r}},$ (2)
$r>0,\ i=1,\dots,n.$
For the rest of this section, we will be assuming that nominal outliers are highly infrequent observations with respect to an underlying discrete distribution. Hence, our arguments will involve Expression (1); extending these to the case of highly frequent itemsets is straightforward.
The formulation in Expression (1) begins by looking at itemsets of unit length and compares their support to a lower threshold value $\sigma_{d}$. If the support lies outside $(\sigma_{d},n]$, the score of all points possessing that itemset is augmented in a way that is directly proportional to $\sigma_{d}$ and inversely proportional to the product of the itemset support and the itemset length raised to a non-negative power $r$. One can immediately notice the similarity of Expression (1) (with $r=1$ and a numerator of a unit) to the one used by [25] and [24]. The idea behind modifying the score to compute $\sigma_{d}/\operatorname{supp}(d)$ instead of just $1/\operatorname{supp}(d)$ is that we wish to ensure a more fair contribution of itemsets with respect to their expected probabilities. For instance, let us suppose that two highly infrequent itemsets $d_{1},d_{2}$ of unit length involving two distinct nominal variables, occur the same number of $\operatorname{supp}(d)$ times in a data set of $n$ observations, respectively. The nominal score of [24] would be the same for observations possessing $d_{1}$ and $d_{2}$. However, if the associated probabilities are given by $\pi_{d_{1}}<\pi_{d_{2}}$, one would expect the relevant threshold values to be $\sigma_{d_{1}}<\sigma_{d_{2}}$. Hence, it would make more sense to treat $d_{2}$ as more outlying compared to $d_{1}$, since the former was expected to occur much more frequently than the latter. Computing the ratio $\sigma_{d}/\operatorname{supp}(d)$ is a way to account for this and ensure that the score is augmented proportionally to the ratio of the minimum expected over the actual support of each highly infrequent itemset.
Our main contributions include a method for determining the support threshold values $\sigma_{d}$ that are specific to each itemset $d$. Moreover, we propose a suitable stopping criterion, that is a $\operatorname*{\operatorname{MAXLEN}}$ value, that accounts for the sparsity introduced in higher-dimensional contingency tables. We also derive some mathematical properties of $s(\cdot)$, such as when $s(\cdot)$ is maximised and what its maximum value is; these are included in Propositions 1 & 2 in Section 4.
3.2 Choice of hyperparameter values
The hyperparameters that are involved in the calculation of the score $s(\cdot)$ are the minimum support threshold $\sigma_{d}$, the maximum length of sequences considered $\operatorname*{\operatorname{MAXLEN}}$ and the exponent term $r$. We start by looking at the minimum support threshold value $\sigma_{d}$. This is defined with respect to the assumed underlying Multinomial distribution for each itemset. Assume without loss of generality that we are working with itemsets $d$ consisting of the first $\lvert d\rvert$ nominal variables. For such itemsets, the underlying Multinomial distribution that we consider is:
$\boldsymbol{X}=\boldsymbol{x}\mid d\sim\text{Multinomial}\left(1,\boldsymbol{p}^{d}\vcentcolon=\bigotimes\limits_{i=1}^{\lvert d\rvert}\boldsymbol{\pi}_{i}\right),$
where $\otimes$ is the Kronecker product. Notice how this formulation implicitly assumes independence of the nominal variables; a dependence structure between the nominal variables would not really change anything besides the definition of the probability vector, which would have to be explicitly provided by the user instead.
Returning to our original problem of determining a suitable threshold value for the support of an itemset $d$, we compute the support of the itemset and see if it falls within a specific range of reasonable values under the assumed model. This is done by constructing an upper one-sided $100(1-\alpha)\%$ confidence interval for the probability of observing itemset $d$. Since all itemsets of length $\lvert d\rvert$ form a Multinomial distribution with $\prod_{i=1}^{\lvert d\rvert}\ell_{i}$ possible outcomes, we use the method of [41] to construct $100(1-\alpha)\%$ simultaneous confidence intervals for the Multinomial proportions. The confidence intervals are constructed by approximating the Multinomial cumulative distribution function (CDF) using a theorem from [31]. The theorem approximates the Multinomial CDF as the product of Poisson probabilities multiplied by the probability mass function of a sum of truncated Poisson random variables, which is approximated using Edgeworth series. The two-sided $100(1-\alpha)\%$ confidence intervals for the components of the Multinomial proportions vector $\boldsymbol{p}^{d}$ are given by:
$\left(p^{d}_{i}-\frac{c_{\alpha}}{n},p^{d}_{i}+\frac{c_{\alpha}+2\gamma_{\alpha}}{n}\right),\quad i=1,\ldots,\prod_{j=1}^{\lvert d\rvert}\ell_{j},$ (3)
where $\gamma_{\alpha}=(1-\alpha-\nu(c_{\alpha}))/(\nu(c_{\alpha}+1)-\nu(c_{\alpha}))$, $\nu(c_{\alpha})=\mathbb{P}(x_{i}-c_{\alpha}\leq X_{i}\leq x_{i}+c_{\alpha})$ with $X_{i}\sim\text{Multinomial}(n,p^{d}_{i})$ and $c_{\alpha}\in\mathbb{Z}$ satisfies $\nu(c_{\alpha})<1-\alpha<\nu(c_{\alpha}+1)$.
In a similar manner, the upper and lower one-sided $100(1-\alpha)\%$ confidence intervals are given by:
$\left(p^{d}_{i}-\frac{c_{2\alpha}}{n},1\right),$ $\quad i=1,\ldots,\prod_{j=1}^{\lvert d\rvert}\ell_{j},$
$\left(0,p^{d}_{i}+\frac{c_{2\alpha}+2\gamma_{2\alpha}}{n}\right),$ $\quad i=1,\ldots,\prod_{j=1}^{\lvert d\rvert}\ell_{j}.$
The confidence intervals obtained for the Multinomial proportions are used to determine what the least expected values for the itemset supports should be. Multiplying the lower bound of each upper one-sided interval by the sample size $n$ yields bounding values for the support of each itemset. These are in fact our values for $\sigma_{d}$, determining when an itemset is highly infrequent with respect to the underlying Multinomial distribution. This extends the idea of [36] and [24], who also augmented the nominal outlyingness scores of observations possessing highly infrequent itemsets using an expression similar to (1), but with a constant $\sigma$ value that was the same for all itemsets of any length.
As the itemset length grows, the associated contingency tables become more sparse. This leads to a serious issue for our procedure if one assumes that $p_{i}^{d}\overset{\lvert d\rvert\rightarrow p}{\longrightarrow}0$, since almost any itemset will be considered as highly infrequent. Besides, allowing for $\lvert d\rvert\rightarrow p$ leads to a substantial and potentially prohibitive increase in computational resources required. As a result, we seek to find a maximum allowed value for $\lvert d\rvert$, which we denote by $\operatorname*{\operatorname{MAXLEN}}$, so that itemsets of length up to $\operatorname*{\operatorname{MAXLEN}}$ will be included in our score computations. Determining $\operatorname*{\operatorname{MAXLEN}}$ so that it reflects on our concerns regarding sparsity can be done by requiring all upper threshold values $\sigma_{d}$ to be at least equal to two for all itemsets $d$. The rationale behind this is that if $\sigma_{d}<2\ \forall d:\lvert d\rvert=M$, then only itemsets of length $M\leq p$ with unit support will not be flagged as infrequent. This is a very restrictive case as most (if not all) observations will have their scores augmented due to being more frequent than expected. Thus, we define $\operatorname*{\operatorname{MAXLEN}}$ as follows:
[…]
$\left.\mathrm{for\ all}\ d\subseteq\boldsymbol{x}_{i}\ \mathrm{with}\ \lvert d\rvert\leq M\right\}.$
The use of concepts from the association rule mining literature for the proposed score of nominal outlyingness presents a practical issue related to computational complexity. For a given value of $\operatorname*{\operatorname{MAXLEN}}$, the total number of contingency tables that can be generated is given by $\sum_{k=1}^{\operatorname*{\operatorname{MAXLEN}}}C^{p}_{k}$, where $C$ denotes the binomial coefficient. Moreover, the number of possible itemsets also gets very large which shows that under the presence of many variables, the total number of comparisons that need to be made can grow very rapidly, especially for large $\operatorname*{\operatorname{MAXLEN}}$. As a result, any computations that may be redundant should be omitted. Hence, we implement support-based pruning, excluding supersets of any highly infrequent itmsets from subsequent computations. That is, if $\lvert d_{1}\rvert<\operatorname*{\operatorname{MAXLEN}}$ and $\operatorname{supp}(d_{1})<\sigma_{d_{1}}$, then any itemset $d_{2}\supset d_{1}$ will not be checked in subsequent calculations.
Support-based pruning, as also implemented in [24], is very convenient when the score is computed in an ascending order from $\lvert d\rvert=1$ up to $\operatorname*{\operatorname{MAXLEN}}$. However, this is only useful in the case of looking for highly infrequent itemsets. In contrast, using the Apriori principle per se is much more useful when we start looking at itemsets of length $\operatorname*{\operatorname{MAXLEN}}$ until we reach individual categorical levels; this approach is favoured when highly frequent itemsets are of interest instead. In order to illustrate this, we present a schematic representation of how the calculation of the score would proceed for a data point $\boldsymbol{x}$ consisting of three nominal levels, say $\boldsymbol{x}=(1,1,1)^{\intercal}$. This is depicted in Figure 1, where we use a tree representation to illustrate how the search for highly frequent or infrequent itemsets is being conducted.
(a) Bottom-up search for highly infrequent itemsets.
(b) Top-down search for highly frequent itemsets.
Figure 1: Implementation of the Apriori principle in support-based pruning for a toy example using a tree representation. Notation $(\boldsymbol{X}_{i},l)$ means $\boldsymbol{X}_{i}=l$. Red and green nodes represent highly infrequent and highly frequent itemsets, respectively.
Figure 1 illustrates how the algorithm proceeds if the user is interested in highly infrequent or highly frequent itemsets. A bottom-up search is being used in the first case; this bottom-up traversal of the tree allows for highly infrequent itemsets to be detected and any of their supersets to be removed from subsequent searching. For instance, in Figure 1(a) the itemset consisting of $\boldsymbol{X}_{1}=1$ and $\boldsymbol{X}_{2}=1$ is infrequent, thus the itemset that further includes the value of $\boldsymbol{X}_{3}$ (no matter what that may be) will not be considered in the search. If highly frequent itemsets are of interest instead, the method will proceed using a top-down approach instead, as shown in Figure 1(b). In this case, if the itemset consisting of $\boldsymbol{X}_{1}=1$ and $\boldsymbol{X}_{3}=1$ is flagged as frequent, all its subsets will also be marked as such and will therefore not be included in the rest of the search.
Our proposal for the score includes an additional hyperparameter, that is the exponent term $r>0$. This term controls the influence of itemsets of larger length in the computation of the score of nominal outlyingness. As [24] pointed out, an observation with frequent itemsets of unit length and $k\leq p$ infrequent itemsets of length two should not be seen as suspicious to be an outlier as an observation with $k$ infrequent itemsets of unit length. Our proposed framework extends this idea by letting the user decide the score contribution of itemsets of greater length. A larger $r$ value will yield a smaller augmentation of $s(\cdot)$ as $\lvert d\rvert$ increases, hence providing additional flexibility and control of how the score of outlyingness is being computed. We cannot recommend a specific value for $r$ as that is typically context-specific and depends on the weight the user wants to assign to highly infrequent (or frequent) itemsets of greater length. For instance, a reasonable value for $r$ would be $r\leq 3$; in case we set $r=3$, we would need $\sigma_{d}/\operatorname{supp}(d)>8$ for a two-dimensional itemset $d$ to contribute to $s(\cdot)$ as much as a specific outlying level on its own. While setting $r>3$ is certainly possible, we suggest that $r$ should not be much larger than the recommended value of $r=3$, so that longer sequences can also contribute to the score, at least to some extent.
3.3 Nominal outlyingness depth & Contribution matrix
We introduce the concept of nominal outlyingness depth, as an extension of our proposed method for quantifying nominal outlyingness. This is an additional metric that we introduce for meta-analysis purposes, so that the scores are easier for the user to understand and interpret. As we have illustrated in Figure 1, the way the score $s(\cdot)$ is being computed is a hierarchical process and can therefore be represented using a tree-based structure. This has motivated us to come up with a concept inspired from the continuous anomaly detection literature and more precisely from the Isolation Forest algorithm [33]. In the Isolation Forest algorithm, the outlyingness of each observation can be assessed by its average depth; that is the average number of random splits that it takes for the observation to be isolated and appear in a region that includes no other points. We proceed in a similar manner and introduce a Nominal Outlyingness Depth concept (NOD), which we denote by $\operatorname{NOD}(\cdot)$, given by:
$\operatorname{NOD}(\boldsymbol{x}_{i})=\left.\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\lvert d\rvert\middle/\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\right.1,$ (4)
$i=1,\dots,n.$
The rationale behind Expression (4) is that for each observation $\boldsymbol{x}_{i}$ we compute the average itemset length over all infrequent itemsets included in $\boldsymbol{x}_{i}$. A similar Expression to (4) can be computed if highly frequent itemsets are of interest; the only modification is changing the interval $(\sigma_{d},n]$ to $[0,\sigma_{d})$ and $\lvert d\rvert$ on the numerator becomes $\operatorname*{\operatorname{MAXLEN}}-\lvert d\rvert+1$, to account for the fact that we use a top-down search instead. One can also consider the definition of the nominal outlyingness depth using a tree representation; that would be the average traversal depth of highly infrequent or highly frequent itemsets, where the depth is relative to a base level that depends on whether a bottom-up or a top-down search is being used. Such an illustration is given in Figure 2 for the toy example of $\boldsymbol{x}=(1,1,1)^{\intercal}$. For instance, according to Figure 2(b), we are looking for highly infrequent itemsets and thus we use a bottom-up approach. The nominal outlyingness depth in this case is given by $(1+3)/2=2$. Similarly for the case of highly frequent itemsets, where a top-down search is being employed as in Figure 2(b), the nominal outlyingness depth would be $(2+3)/2=5/2$. The depth value ranges from a unit up to $\operatorname*{\operatorname{MAXLEN}}$, with a lower value corresponding to a larger degree of outlyingness, except for observations with no highly frequent or highly infrequent itemsets; these have a zero depth.
(a) Nominal outlyingness depth calculation for highly infrequent itemsets.
(b) Nominal outlyingness depth calculation for highly frequent itemsets.
Figure 2: Illustration of the calculation of nominal outlyingness depth calculation using a tree representation. Notation $(\boldsymbol{X}_{i},l)$ means $\boldsymbol{X}_{i}=l$. Red and green nodes represent highly infrequent and highly frequent itemsets, respectively. Support-based pruning has been applied in both cases.
The final concept we introduce is that of the contribution matrix, which we denote by $\boldsymbol{\mathrm{C}}$. This is a $(n\times p)$-dimensional matrix, with the $(i,j)$th entry being the contribution of variable $j$ to the nominal outlyingness score of $\boldsymbol{x}_{i}$. The contribution of the $j$th variable to the nominal score of outlyingness of the $i$th observation for highly infrequent itemsets is given by:
$c_{i,j}=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ j\subseteq d\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r+1}},\quad r>0,$ (5)
$i=1,\dots,n,\ j=1,\dots,p.$
As a result, all the elements of the contribution matrix will be non-negative and the sum of each row is in fact the score of nominal outlyingness for the relevant observation, that is $\sum_{j=1}^{p}c_{i,j}=s(\boldsymbol{x}_{i})$. The main idea behind Expression (5) is that if an itemset is found to be highly infrequent, all variables involved will get an equal contribution to the augmentation of the score that is caused by this sequence of categorical levels. One can therefore use the contribution matrix to facilitate understanding of the source of nominal outlyingness of the data by seeing which variables are involved in highly infrequent (or frequent) itemsets, causing an increase of the score of nominal outlyingness.
4 Properties of the proposed score
We now explore some properties of our proposed score of nominal outlyingness. Notice that we focus exclusively on the case of an outlier being defined as consisting of highly infrequent itemsets; extending these properties to the frequent case is straightforward. We are mainly interested in the range of values the score can take, as well as on when it is maximised; essentially, we would like to know under what circumstances an observation is extremely likely to be outlying and how the associated hyperparameters influence its score.
Proposition 1.
The maximum value of the nominal score of outlyingness $s(\boldsymbol{x}_{i})$ for infrequent itemsets in a data set with $p$ nominal variables is attained for all itemsets of length $1\leq k\leq p$ appearing just once.
Proof.
We assume that we have $p\geq 2$ nominal variables (for the case of $p=1$, the result follows immediately by the definition of the nominal score) and recall the formulation of the discrete score for an observation $1\leq i\leq n$ as given in Expression (1):
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r}},\quad r>0.$
Assuming a large $\operatorname*{\operatorname{MAXLEN}}$ value (to avoid having many restrictions on the calculation of the score), it is easy to see that itemsets of unit support with a large minimum support threshold value $\sigma_{d}$ yield a higher increase in the score. This case corresponds to highly misspecified large probabilities for itemsets that only appear once in the data, for which $\sigma_{d}=n-\delta$, where $\delta$ is typically a small positive integer. Since $\sigma_{d}$ decreases as $\lvert d\rvert$ gets larger and as this is a parameter that is hard to determine in advance, we will stick with the maximisation of the $1/\lvert d\rvert^{r}$ term, assuming the maximum possible $\sigma_{d}$ value of $\sigma_{d}=n-1$. Therefore, the total score becomes equal to the following expression:
$\mathcal{A}=(n-1)\times\\ \times\left\{p-\sum\limits_{i=1}^{p-1}\alpha_{i}+\sum\limits_{i=1}^{p-1}\dbinom{\alpha_{i}}{i+1}\frac{1}{\left(i+1\right)^{r}}\right\},\\ \alpha_{i}\in\left\{0,i+1,\dots,p\right\},\ \sum\limits_{i=1}^{p-1}\alpha_{i}\leq p.$
Notice that the first two terms inside the curly brackets correspond to the contribution of the itemsets of unit length that appear once in the data set, while the third term corresponds to the contribution of all possible itemsets of greater length (up to length $p$) to the score. The coefficients $\alpha_{i}$ represent the number of nominal variables which are included in infrequent itemsets (of unit support) of length $i+1$. We further impose the restriction that the sum of the $\alpha_{i}$’s is at most equal to $p$, since any itemsets of length $i+1$ are defined based on $i+1$ variables, and we are restricted to $p$ nominal features. Moreover, we could potentially have no infrequent itemsets of length $i+1$ (in which case $\alpha_{i}=0$) but if we do have any, then these should be observed in at least $i+1$ nominal features. Despite the abuse of notation, we assume that $C^{0}_{i+1}=0$ for convenience.
Our goal is to find the set of values $\boldsymbol{\alpha}=\left(\alpha_{1},\dots,\alpha_{p-1}\right)$ that maximise $\mathcal{A}$ and more precisely, we aim to show that this is achieved either when all the $\alpha_{i}$’s are equal to zero or when we are on the boundary of the solution space, i.e. when all the $\alpha_{i}$’s sum to $p$ and all of them but one are exactly zero. The first case corresponds to maximising the nominal score for all $p$ nominal features of an observation being unique within the data set, while the second is interpreted as achieving the maximum nominal score possible when all itemsets of length $i+1$ that can be generated from $p$ nominal variables appear just once and there exist no itemsets of greater or lower length which are infrequent. In both these cases, we basically end up with the conclusion that the maximum score is attained for all itemsets of one specific length occurring once in the data set.
For $p<2^{r+1}+1$, we would rather have $\alpha_{i}=0\ \forall i$ if:
$\sum\limits_{i=1}^{p-1}\alpha_{i}>\sum\limits_{i=1}^{p-1}\dbinom{\alpha_{i}}{i+1}\frac{1}{(i+1)^{r}}.$
The above condition can be equivalently written, by some straightforward algebraic manipulations, as:
$\sum\limits_{i=1}^{p-1}\left\{\alpha_{i}\times\left(\frac{\left(\alpha_{i}-1\right)\times\dots\times\left(\alpha_{i}-i\right)}{(i+1)^{r+1}\times i!}-1\right)\right\}<0.$
Since we assume that all the $\alpha_{i}$’s are non-negative, the expression above can be negative if and only if $\exists\alpha_{i}>0$ such that:
$F$ $=\frac{\left(\alpha_{i}-1\right)\times\dots\times\left(\alpha_{i}-i\right)}{(i+1)^{r+1}\times i!}-1$ (6)
$=\frac{\Gamma\left(\alpha_{i}\right)}{\Gamma\left(\alpha_{i}-i\right)\times\Gamma\left(i+1\right)\times(i+1)^{r+1}}-1<0,$
where $\Gamma(\cdot)$ is the gamma function. If $F<0\ \forall\alpha_{i}>0$, we can ensure that $\mathcal{A}$ is maximised for all the $\alpha_{i}$’s being equal to zero. Since Expression (6) is maximised when $i=1$ and $\alpha_{1}=p$, we obtain:
$F=\frac{p-1}{2^{r+1}}<1\iff p<2^{r+1}+1.$
Therefore, $\alpha_{i}=0\ \forall i$ maximises expression $\mathcal{A}$ for $p<2^{r+1}+1$.
We now assume $p\geq 2^{r+1}+1$. We first show that getting to the solution boundary is always preferable when having one non-zero $\alpha_{i}$ and then, we show that setting $\alpha_{j}$ equal to a non-zero value, where $i\neq j$, decreases the value of $\mathcal{A}$. We begin by assuming that there exists an index $j$ such that $\alpha_{j}>0$ and $\alpha_{i}=0\ \forall i\neq j$ and we look at what happens to the value of $\mathcal{A}$ if we decrease $\alpha_{j}$ by a unit. Due to the restriction on the values that $\alpha_{j}$ can take, we need to assume that $j\neq p-1$, since $\alpha_{p-1}$ can only be equal to 0 or $p$. We denote the loss that results from decreasing $j$ by a unit by $G_{1}$ and that is equal to:
$G_{1}$ $=p-\alpha_{j}+\dbinom{\alpha_{j}}{j+1}\times\frac{1}{(j+1)^{r}}-$
$-\left\{p-\alpha_{j}+1+\dbinom{\alpha_{j}-1}{j+1}\times\frac{1}{(j+1)^{r}}\right\}$
$=\frac{1}{(j+1)^{r}}\left\{\dbinom{\alpha_{j}}{j+1}-\dbinom{\alpha_{j}-1}{j+1}\right\}-1$
$=\frac{\Gamma\left(\alpha_{j}\right)}{(j+1)^{r-1}\times\Gamma\left(j+2\right)\times\Gamma\left(\alpha_{j}-j\right)}-1.$
When decreasing $\alpha_{j}$ from its minimum possible non-zero value of $j+1$ to zero, we define $G_{2}$ to be the loss corresponding to this decrease:
$G_{2}$ $=p-\alpha_{j}+\dbinom{\alpha_{j}}{j+1}\times\frac{1}{(j+1)^{r}}-p$
$=-(j+1)+\dbinom{j+1}{j+1}\times\frac{1}{(j+1)^{r}}$
$=\frac{1}{(j+1)^{r}}-(j+1).$
Hence, for $G<0\ \forall j\in\mathbb{Z}^{+}$, which implies $\mathcal{A}$ increases when going from $\alpha_{j}=j+1$ to $\alpha_{j}=0$. Recalling that $\alpha_{j}>j$, we may write $\alpha_{j}=j+z$, where $z\in\{1,\ldots,p-j\}$. The expression for $G_{1}$ then becomes:
$G_{1}=\frac{\Gamma\left(j+z\right)}{(j+1)^{r-1}\times\Gamma\left(j+2\right)\times\Gamma\left(z\right)}-1.$
The above expression is increasing in $z$ (this can be seen by considering the ratio $\Gamma(j+z)/\Gamma(z)$ and using Stirling’s approximation for the gamma function) and decreasing in $j$. Thus $G_{1}$ is minimised when $j=p-2$ and $z=1$, in which case:
$G_{1}$ $=\frac{\Gamma\left(p-1\right)}{(p-1)^{r-1}\times\Gamma\left(p\right)\times\Gamma\left(1\right)}-1$
$=\frac{1}{(p-1)^{r}}-1<0,$
since $p\geq 2^{r+1}+1>2$ (recall $r>0$). Thus, when we are on the boundary of the solution space, it is preferred for us to stay there, while if we are at the minimum non-zero value of $\alpha_{j}$, it is preferred to simply set it to zero.
The final part of the proof consists of showing that if we are on the boundary of the solution space, with $\alpha_{j}=p$ and $\alpha_{i}=0\ \forall i\neq j$, then activating $\alpha_{k}$ so that $\alpha_{k}=k+1$ (where $k\neq j$) leads to a decrease in the value of $\mathcal{A}$. Now clearly, if $\alpha_{k}$ is set equal to $k+1$, the value of $\alpha_{j}$ has to drop to $p-(k+1)$, so that the boundary constraint is not violated. We define the loss incurred by activating $\alpha_{k}$ by $H$:
$H$ $=\dbinom{p}{j+1}\times\frac{1}{(j+1)^{r}}-$
$-\left\{\dbinom{p-k-1}{j+1}\frac{1}{(j+1)^{r}}+\dbinom{k+1}{k+1}\frac{1}{(k+1)^{r}}\right\}$
$=\frac{1}{(j+1)^{r}}\times\left\{\dbinom{p}{j+1}-\dbinom{p-k-1}{j+1}\right\}-$
$-\frac{1}{(k+1)^{r}}$
$=\frac{1}{(j+1)^{r}\times\Gamma\left(j+2\right)}\times$
×$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-j\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-k-j-1\right)}\right\}-\frac{1}{(k+1)^{r}}.$
We distinguish between two possible cases, namely $j<k$ and $j>k$. In the former case, we obtain:
$H$ $=\frac{1}{(k-\beta+1)^{r}\times\Gamma\left(k-\beta+2\right)}\times$
×$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-k+\beta\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-2k+\beta-1\right)}\right\}-$
$-\frac{1}{(k+1)^{r}},$
while in the latter:
$H$ $=\frac{1}{(k+\beta+1)^{r}\times\Gamma\left(k+\beta+2\right)}\times$
×$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-k-\beta\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-2k-\beta-1\right)}\right\}-$
$-\frac{1}{(k+1)^{r}},$
where $\beta\in\mathbb{Z}^{+}$. These expressions for $H$ give us additional restrictions, which are that $p-2k+\beta-1>0$ and $p-2k-\beta-1>0$, so that all terms are well-defined. We now show that $H>0\ \forall k\neq j$ just for the case of $j<k$; the second case follows analogously. Let the first term of $H$ be denoted by $H_{1}$ and denote the second one by $H_{2}$. It is obvious that $\beta=1$ minimises $H_{1}$. Thus, we have:
$H$ $=\frac{1}{k^{r}\times\Gamma(k+1)}\left\{\frac{\Gamma(p+1)}{\Gamma(p-k)}-\frac{\Gamma(p-k)}{\Gamma(p-2k)}\right\}-$
$-\frac{1}{(k+1)^{r}}.$
Now we can write the ratio of $H_{1}$ over $H_{2}$ as:
$\frac{H_{1}}{H_{2}}=\frac{(k+1)^{r}\times Q(k)}{k^{r}\times\Gamma(k+1)}=\frac{(1+1/k)^{r}\times Q(k)}{\Gamma(k+1)},$
where $Q(k)=\Gamma(p+1)/\Gamma(p-k)-\Gamma(p-k)/\Gamma(p-2k)$. It is also easy to see that $Q(k)$ is increasing for $k$, therefore its minimum is attained when $k=2$. In that case, we have:
$\frac{H_{1}}{H_{2}}$ $=\frac{(k+1)^{r}\times Q(k)}{k^{r}\times\Gamma(k+1)}$
$=\frac{(1+1/k)^{r}\times Q(k)}{\Gamma(k+1)}$
$\geq\frac{(3/2)^{r}\times Q(2)}{2}>1.$
Thus, the minimum value of $H<1$ which shows $H>0\ \forall k\neq j$. As a result activating $\alpha_{k}$ is not preferred as it leads to a positive loss value.
Thus, we have shown the following:
  1. 1.
    The greatest $\mathcal{A}$ value for $p<2^{r+1}+1$ is achieved when all the $\alpha_{i}$’s are equal to zero.
  2. 2.
    When $p\geq 2^{r+1}+1$, $\mathcal{A}$ is maximised when we are on the boundary of the solution space.
  3. 3.
    If $p\geq 2^{r+1}+1$ and we are on the boundary of the solution space with just one parameter $\alpha_{j}$ being equal to $p$, then activating any other parameter $\alpha_{k}$ (where $j\neq k$) so that it is not longer equal to zero, leads to a lower value of $\mathcal{A}$.
Given the above, we can conclude that the optimal solution for $p<2^{r+1}+1$ is attained when all parameters are equal to zero, while for $p\geq 2^{r+1}+1$, the optimal solution is on the boundary of the solution space with just one parameter $\alpha_{i}$ being non-zero and thus equal to $p$, as required. ∎
Given Proposition 1, we can now proceed to find what the maximum possible score of nominal outlyingness is. This is derived in Proposition 2.
Proposition 2.
The maximum value of the score of nominal outlyingness $s(\boldsymbol{x}_{i})$ for the $i$th observation of a data set with $p$ nominal variables is given by $p(n-1)$ for $p<2^{r+1}+1$, and by the following expression for $p\geq 2^{r+1}+1$:
$s^{\mathrm{max}}(\boldsymbol{x}_{i})=(n-1)\frac{\dbinom{p}{\min\left\{\operatorname*{\operatorname{MAXLEN}},\left\lfloor\frac{p-r}{2}\right\rfloor\right\}}}{\left(\min\left\{\operatorname*{\operatorname{MAXLEN}},\left\lfloor\frac{p-r}{2}\right\rfloor\right\}\right)^{r}},$
[…]
Proof.
As proven in Proposition 1, if we assume that the largest possible value of $\sigma_{d}/\operatorname{supp}(d)$ is given by $n-1$, the score of nominal outlyingness $s$ is maximised when we only have highly infrequent (or frequent) itemsets of unit length, for $p<2^{r+1}+1$. In this case, it is obvious that the total number of such itemsets is equal to $p$, thus the score becomes equal to $(n-1)p$.
Now for the case of $p\geq 2^{r+1}+1$, as previously shown, the largest possible score can be achieved on the boundary of the solution space; that is for all itemsets of a specific length $k\leq p$ having unit support. The total score is then given by $(n-1)\times C^{p}_{k}/{k^{2}}$,where $C$ is the binomial coefficient. We therefore seek to find the itemset length $k^{*}\in\mathbb{Z}^{+}$ that maximises this expression of the score, for fixed $p$. This means that the score for $k=k^{*}-1$ and for $k^{*}+1$ (we assume $1<k^{*}<p$) will be less than that for $k^{*}$, or equivalently (omitting the common $(n-1)$ term):
$\dbinom{p}{k^{*}+1}\frac{1}{\left(k^{*}+1\right)^{r}}-\dbinom{p}{k^{*}}\frac{1}{\left(k^{*}\right)^{r}}$ $<0,$ (7)
$\dbinom{p}{k^{*}}\frac{1}{\left(k^{*}\right)^{r}}-\dbinom{p}{k^{*}-1}\frac{1}{\left(k^{*}-1\right)^{r}}$ $>0.$ (8)
Starting with Expression (7), we have:
$\dbinom{p}{k^{*}+1}\frac{1}{\left(k^{*}+1\right)^{r}}-\dbinom{p}{k^{*}}\frac{1}{\left(k^{*}\right)^{r}}=\\ =\frac{p!}{\left(k^{*}+1\right)!\left(p-k^{*}-1\right)!}\frac{1}{\left(k^{*}+1\right)^{r}}-\\ -\frac{p!}{\left(k^{*}\right)!\left(p-k^{*}\right)!}\frac{1}{\left(k^{*}\right)^{r}}\\ =\frac{p!}{\left(k^{*}\right)!\left(p-k^{*}-1\right)!\left(k^{*}\right)^{r}}\times\\ \times\left\{\frac{1}{\left(k^{*}+1\right)\left(1+\frac{1}{k^{*}}\right)^{r}}-\frac{1}{p-k^{*}}\right\},$
and since we know that the first term is strictly positive, it suffices to show that the expression inside the curly brackets is negative. We bring this to the following form:
[…]
hence it suffices to show that the numerator is negative, since the denominator is again strictly positive. This gives:
$\left(p-k^{*}\right)-\left(k^{*}+1\right)\left(1+\frac{1}{k^{*}}\right)^{r}<0$
$\iff$ $p<k^{*}+\left(k^{*}+1\right)\left(1+\frac{1}{k^{*}}\right)^{r}.$ (9)
We proceed similarly to derive an additional bound based on Expression (8). More precisely:
$\dbinom{p}{k^{*}}\frac{1}{\left(k^{*}\right)^{r}}-\dbinom{p}{k^{*}-1}\frac{1}{\left(k^{*}-1\right)^{r}}=\\ =\frac{p!}{\left(k^{*}\right)!\left(p-k^{*}\right)!}\frac{1}{\left(k^{*}\right)^{r}}-\\ -\frac{p!}{\left(k^{*}-1\right)!\left(p-k^{*}+1\right)!}\frac{1}{\left(k^{*}-1\right)^{r}}\\ =-\frac{p!}{\left(k^{*}-1\right)!\left(p-k^{*}\right)!\left(k^{*}-1\right)^{r}}\times\\ \times\left\{\frac{1}{p-k^{*}+1}-\frac{\left(1-\frac{1}{k^{*}}\right)^{r}}{k^{*}}\right\},$
so by noticing that the first term is strictly positive and the product should be positive, we require that the expression inside the curly brackets is negative. A bit of re-arrangement gives:
$\frac{1}{p-k^{*}+1}-\frac{\left(1-\frac{1}{k^{*}}\right)^{r}}{k^{*}}=$
$\frac{k^{*}-\left(p-k^{*}+1\right)\left(1-\frac{1}{k^{*}}\right)^{r}}{\left(p-k^{*}+1\right)k^{*}},$
thus we require that the numerator is negative, since the denominator is strictly positive. Hence, we have:
$k^{*}-\left(p-k^{*}+1\right)\left(1-\frac{1}{k^{*}}\right)^{r}<0$
$\iff$ $p>\frac{k^{*}}{\left(1-\frac{1}{k^{*}}\right)^{r}}+k^{*}-1.$ (10)
[…]
Figure 3: Plot of bounds of $p$ for varying $r$. The blue shaded region corresponds to the upper bound from Expression (9) and the red shaded region corresponds to the lower bound from Expression (10). The green shaded region is the region where both these bounds are satisfied. The vertical dashed line at $k=1$ is the asymptote of the expression for the lower bound. The yellow dotted line is given by $p=2k+r$; this is the line approximating the integer solutions in the green shaded region as $p\rightarrow\infty$.
Figure 4: Plot of bounds of $p$ for $r=2$. The rectangular region $[6,8]\times[10,24]$ on the left subplot is zoomed in on the right, illustrating how the 2 boundary curves become parallel as $p$ increases.
The points of intersection of the boundary curves is (as expected) given by $p=2^{r+1}+1$ and solving one of Expressions (9) or (10) for the corresponding $k$ value. We can see from both Figures 3 & 4 that the two boundary curves become parallel as $k\rightarrow+\infty$. We can compute their gradient to validate this:
$\frac{\mathrm{d}\left(k+\left(k+1\right)\left(1+\frac{1}{k}\right)^{r}\right)}{\mathrm{d}k}=\\ =1+\left(1+\frac{1}{k}\right)^{r}-r\frac{k+1}{k^{2}}\left(1+\frac{1}{k}\right)^{r-1}\overset{k\rightarrow\infty}{\longrightarrow}2,\\ \frac{\mathrm{d}\left(\frac{k}{\left(1-\frac{1}{k}\right)^{r}}+k-1\right)}{\mathrm{d}k}=\\ =1+\left(1-\frac{1}{k}\right)^{-r}-\frac{r}{k}\left(1-\frac{1}{k}\right)^{-r-1}\overset{k\rightarrow\infty}{\longrightarrow}2.$
Therefore, we can access the set of solutions (corresponding to the region where both bounds are satisfied) by considering the line $p=2k+\alpha$. The constant $\alpha$ can be computed by noticing that this line will pass through the midpoint of each line segment defined by the two boundary curves for large $k$. Hence, we need to compute the difference between the two boundary curves:
$k+\left(k+1\right)\left(1+\frac{1}{k}\right)^{r}-\frac{k}{\left(1-\frac{1}{k}\right)^{r}}+k-1=$
$=\frac{(k+1)(k^{2}-1)^{r}-k^{2r+1}}{k^{r}(k-1)^{r}}+1\overset{k\rightarrow\infty}{\longrightarrow}2,$
since the first term in the numerator involves a $k$ term of degree $12r+$ 1 (which cancels with the negative term on the numerator) and a term $k^{2r}$ that simplifies to $1$ upon division. The rest of the terms tend to 0 as $k\rightarrow\infty$ so the only term that remains is the unit coefficient of $k^{2r}$, which means that for large $k$ the vertical distance between the two boundary curves is equal to two units. Their midpoint is just one unit away from each curve, thus we require:
$k+(k+1)\left(1+\frac{1}{k}\right)^{r}-2k-\alpha\overset{k\rightarrow\infty}{\longrightarrow}1,$
which is satisfied for $\alpha=r$. This is shown below using the binomial series expansion of $(1+1/k)^{r}$:
$k+(k+1)\left(1+\frac{1}{k}\right)^{r}-2k-\alpha=$
$=k+(k+1)\left(1+\frac{r}{k}+\mathcal{O}(k^{-2})\right)-2k-\alpha$
$=k+k+r+1+\mathcal{O}(k^{-1})-2k-\alpha$
$=r+1-\alpha+\mathcal{O}(k^{-1})\overset{k\rightarrow\infty}{\longrightarrow}1,$
[…]
The proof is completed by considering that $\operatorname*{\operatorname{MAXLEN}}$ is always at least equal to one; thus the maximum score for $p<2^{r+1}+1$ is equal to $p$, while for $p\geq 2^{r+1}+1$, the maximum score is equal to the expression that we mentioned at the beginning of the proof, evaluated at $k=k^{*}$. However, since $\operatorname*{\operatorname{MAXLEN}}$ can be less than $k^{*}$ and given that the maximum score expression is increasing for $k\in\mathbb{Z}^{+}_{\leq k^{*}}$, the maximum score of nominal outlyingness is attained for $k=\min\{\lfloor(p-r)/2\rfloor,\operatorname*{\operatorname{MAXLEN}}\}$. ∎
5 Applications
In this section we illustrate the validity of our proposed method via a simulation study including synthetic data sets. Moreover, we apply our algorithm on four publicly available data sets from the UCI Machine Learning Repository [23] and compare its performance with that of two state-of-the-art algorithms from the frequent pattern mining literature.
5.1 Simulation Study
In order to illustrate the validity of our method, we conduct a simulation study on artificial data. Our proposed framework, which we denote by SONO (Scores Of Nominal Outlyingness) is benchmarked against two frequent pattern mining algorithms, specifically Frequent Pattern Outlier Factor (FPOF) [21] and Frequent Pattern Isolation (FPI) [26]. These methods present certain similarities with SONO in the way they define outliers, yet they require specifying values for $\operatorname*{\operatorname{MAXLEN}}$ and for the minimum relative support threshold. The minimum relative support threshold is analogous to the expression $\sigma_{d}/n$ under our framework, where $n$ denotes the number of observations in the data.
We generate our artificial data set $\boldsymbol{X}$ using the following model:
$\boldsymbol{X}_{1}\sim\text{Categorical}\left(q_{1},\underbrace{\frac{1-q_{1}}{\ell-1},\ldots,\frac{1-q_{1}}{\ell-1}}_{(\ell-1)\text{ times}}\right),$
$\boldsymbol{X}_{2}\sim\text{Categorical}\left(q_{2},\underbrace{\frac{1-q_{2}}{\ell-1},\ldots,\frac{1-q_{2}}{\ell-1}}_{(\ell-1)\text{ times}}\right),$
[…]
In the above, $\boldsymbol{X}_{j}$ denotes the $j$th variable, $p$ is the number of variables and $\ell$ is the number of categorical levels. The last $p-2$ variables are essentially sampled from a discrete uniform distribution, while the first level for the first two variables occurs with probabilities $q_{1}$ and $q_{2}$, respectively. We set $q_{1}=0.10$ and $q_{2}=0.05$ in order to have one level being significantly less frequent than the rest. We vary the values of $n$, $p$ and $\ell$ and allow these to take values $n=200,500,1000$, $p=3,5,7$ and $\ell=2,3,5$. This leads to 27 different scenarios, simulated a hundred times each (using a different seed number), thus yielding 2700 artificial data sets. For each data set, we run FPOF and FPI with a minimum relative support threshold of $1/\ell$ and $\operatorname*{\operatorname{MAXLEN}}$ determined by our method using the relevant functions from the R package fpmoutliers. SONO is run three times; once assuming equal probabilities across all levels for all variables, once with the true probabilities, and once with probability vectors $\left(q_{j}/2,(1-q_{j}/2)/(\ell-1),\ldots,(1-q_{j}/2)/(\ell-1)\right)^{\intercal}$, $j=1,2$, for variables $\boldsymbol{X}_{1},\boldsymbol{X}_{2}$ and equal probabilities across each level for the remaining $p-2$ variables. We record the average rank for the observations possessing the first level for any of $\boldsymbol{X}_{1}$ or $\boldsymbol{X}_{2}$, the estimated value of $\operatorname*{\operatorname{MAXLEN}}$ and the average runtime for each of the three algorithms across twenty-five repetitions. Since some observations may end up with identical scores, we use the highest rank to resolve any ties (a higher rank here refers to a larger score of outlyingness). We also subtract the scores obtained with FPOF from a unit for their ranks to be comparable to these computed using SONO and FPI (FPOF assigns higher scores to inliers). We finally set $r=1$ to be more consistent with FPOF and FPI and $\alpha=0.05$ when running SONO.
In Figure 5 we can see the median rank for observations for which $\boldsymbol{X}_{1}=1$ or $\boldsymbol{X}_{2}=1$, as obtained by FPI, FPOF and SONO for a varying number of variables and observations. For the latter method, the three different cases related to the assumed probabilities are also included. When SONO is given the true probabilities, the median rank is constantly equal to a unit; this comes as no surprise since all scores of outlyingness are equal to zero, thus all observations get the exact same rank of one. Incorporating this contextual information and thus not flagging any observation as potentially outlying is not possible with FPI and FPOF, with both algorithms consistently treating the observations of interest as potential outliers. SONO has also managed to flag these and has even assigned them a higher rank than FPI and FPOF under the assumption of equal level probabilities. On the other hand, when the assumed probabilities for the first levels of $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ are smaller than the true ones, we see that the median rank is much lower and that is because the concept of infrequent itemsets is no longer applicable, further highlighting the validity of our method.
We finally report the median runtime in seconds for our experiments, which heavily depends on the value $\operatorname*{\operatorname{MAXLEN}}$, in Table 2. The reported times are only for the case of probability levels being set equal across each variable to ensure a fair comparison between the three methods. As we can see, SONO is quicker than FPI and FPOF when $\operatorname*{\operatorname{MAXLEN}}=1$ and consistently faster than FPI for $\operatorname*{\operatorname{MAXLEN}}\geq 3$. The median runtime of SONO increases on a large scale as $\operatorname*{\operatorname{MAXLEN}}$ grows (potentially due to its current implementation not being fully optimised for speed), yet it does not take extremely long to produce results which are typically more reliable than these of FPOF and FPI, as seen previously. Finally, we notice that a $\operatorname*{\operatorname{MAXLEN}}$ value of seven was only attained in 181 out of the 900 data sets consisting of seven variables and only when the true underlying level probabilities were provided. This highlights the efficiency of using support-based pruning, leading to a substantially smaller runtime in approximately 80% of the data sets with seven variables.
Figure 5: Median rank of observations for which $\boldsymbol{X}_{1}=1$ or $\boldsymbol{X}_{2}=1$ using FPI, FPOF and SONO with equal probabilities across levels, the true level probabilities and with half the true probabilities for the first levels of $\boldsymbol{X}_{1},\boldsymbol{X}_{2}$.
MAXLEN Algorithm SONO FPOF FPI
1 0.03 0.04 1.18
2 0.67 0.08 1.19
3 0.54 0.08 1.18
4 1.69 0.04 1.02
5 3.75 0.08 1.19
6 24.2 0.14 1.29
Table 1: Median runtime in seconds for SONO, FPI and FPOF for MAXLEN values of one up to six, assuming equal probabilities across variable levels. Smallest times per $\operatorname*{\operatorname{MAXLEN}}$ value are bolded.
5.2 Applications on real data
We implement our method and compare its performance to that of FPOF and FPI on four publicly available data sets from the UCI Machine Learning Repository. The data sets used are the Solar Flare, the Differentiated Thyroid Cancer Recurrence, the Lymphography and the Wisconsin Breast Cancer data sets. These data sets consist of mainly nominal variables and the few variables of a different type (if any) are removed in the cleaning process, so that our method can be applied. Notice that these data sets have previously been used for classification purposes [see for example 34, 15, 9, 5, 8]. The Lymphography and Wisconsin Breast Cancer data sets have also been used for anomaly detection purposes in multiple instances [for instance in 21, 29, 49, 2, 26, among others].
The number of observations for each data set upon cleaning, as well as the number of variables used for each data set are summarised in Table 2, together with the value of $\operatorname*{\operatorname{MAXLEN}}$ as determined by SONO. Notice that we set $\alpha=0.05$ and $r=1$ for SONO, while the probabilities were set to be equal across variable categories. Following the implementations of FPOF and FPI on the Lymphography and Wisconsin Breast Cancer data sets from the papers that introduced these algorithms, we set the minimum relative support threshold to be $0.1$. The value of $\operatorname*{\operatorname{MAXLEN}}$ was set to be the same as the one estimated by SONO for each data set. We also report the proportion of outliers as a percentage for each data set, assuming that instances from the minority class are outlying for the Thyroid Cancer, Breast Cancer and Lymphography data sets. For the Solar Flare data set, observations corresponding to moderate or severe solar flares are taken to be the outliers.
Data set Observations Variables $\operatorname*{\operatorname{MAXLEN}}$ Proportion of Outliers (%)
Solar Flare 1389 10 4 5.26%
Thyroid Cancer 383 13 1 28.20%
Wisconsin Breast Cancer 683 9 1 34.99%
Lymphography 148 18 1 4.05%
Table 2: Number of observations upon data cleaning, number of variables used for nominal outlier detection, $\operatorname*{\operatorname{MAXLEN}}$ values and proportion of outliers for real-world data sets used.
Following the recommendations of [12], we compare the performance of the three algorithms using three evaluation measures. The first one we use is the Detection Rate at Top $K\%$, defined as the proportion of outliers detected within the observations with the top $K\%$ of outlyingness scores. This is analogous to the Recall measure, so we denote this by $\mathcal{R}(K)$. Assuming $O$ is the set of outliers, $\mathcal{R}(K)$ is defined as follows:
$\mathcal{R}(K)=\frac{\lvert\{o\in O:\text{rank}(o)\leq\lceil Kn/100\rceil\}\rvert}{\lvert O\rvert},$
where $\text{rank}(o)$ denotes the rank of an outlying instance $o\in O$ in a sorted list of outlyingness scores (ranked in descending order, with a rank of one being the most outlying). We also compute the average rank for the outlying observations and Area Under the Receiver Operating Characteristic Curve (ROC AUC) value for each method. The ROC AUC value for a list of ranked scores is computed as follows:
ROC AUC $=\frac{1}{\lvert O\rvert\times\lvert I\rvert}\sum_{o\in\lvert O\rvert}\sum_{i\in\lvert I\rvert}\mathcal{S}(o,i),$
$\mathcal{S}(o,i)$ =$=\left\{\begin{array}[]{ll}1,&\text{if rank}(o)>\text{rank}(i)\\ \frac{1}{2},&\text{if rank}(o)=\text{rank}(i)\\ 0,&\text{if rank}(o)<\text{rank}(i)\\ \end{array}\right.,$
where $I$ is the set of inliers [20]. A value closer to one indicates perfect performance, as this implies that most outliers are ranked higher than most of the inliers. In contrast, a value closer to zero will occur when an algorithm assigns higher outlyingness scores to inliers. We report $\mathcal{R}(K)$, the average rank and the ROC AUC for the $K\%$ top-ranked observations for each data set and each algorithm in Table 3 and Figures 6 & 7 for several values of $K$.
Data set SONO FPOF FPI
Solar Flare 377.48 348.34 321.84
Thyroid Cancer 101.74 92.04 99.45
Wisconsin Breast Cancer 116.63 125.67 126.08
Lymphography 4.00 4.50 4.00
Table 3: Average rank of outlying observations in four UCI data sets based on the scores of outlyingness computed using the SONO, FPOF and FPI algorithms. Values are presented in two decimal places and a lower value indicates a higher ranking. Highest average rankings for each data set are bolded.
Figure 6: Detection Rate at Top $K\%$ ($\mathcal{R}(K)$) for observations with the top $K\%$ scores of outlyingness for FPI, FPOF and SONO algorithms on the Solar Flare, Thyroid Cancer, Wisconsin Breast Cancer and Lymphography data sets. The values of $K$ presented are 1, 2.5, 5 and values from 10 up to 50, in steps of 10.
Figure 7: ROC AUC for observations with the top $K\%$ scores of outlyingness for FPI, FPOF and SONO algorithms on the Solar Flare, Thyroid Cancer, Wisconsin Breast Cancer and Lymphography data sets. The values of $K$ presented are 1, 2.5 and values from 5 up to 100, in steps of 5.
Based on the average rank of outliers in Table 3, we see that SONO outperforms FPOF and FPI on the Wisconsin Breast Cancer data, while it achieves better performance than FPOF on the Lymphography data set. Despite not being able to do as well as its competitors on the Solar Flare and the Thyroid Cancer data sets, it has still managed to recover a good amount of outliers, as can be seen in Figure 6. In fact, for the Solar Flare data set, we see that the Detection Rate $\mathcal{R}(K)$ is comparable and even higher than that of FPOF or FPI for $K\leq 30\%$. Similarly for the Thyroid Cancer Data set, the largest scores of outlyingness obtained with each of the three methods correspond in their majority to outliers. Moreover, the ROC AUC for SONO takes very similar values to the ones obtained with FPI and FPOF, as seen in Figure 7. Hence, in a practical application, where the proportion of outliers would be assumed to be less than $50\%$, SONO would be a competitive algorithm to the ones presented, being able to correctly detect over 70% of the outliers.
Overall, SONO has exhibited comparable performance to FPI and FPOF. However, it is important to stress that SONO required far less user input than its competitors, since the value for $\operatorname*{\operatorname{MAXLEN}}$ was estimated from the data and none of the minimum support threshold values was input by the user. Potentially different values for the minimum relative support threshold could have produced much different results for FPI and FPOF, yet these are not explored here and in fact, no recommendations have been made regarding the selection of an optimal value for this hyperparameter. A final appealing feature of SONO is that we can interpret the scores obtained and see which variables have contributed to these using the additional concepts introduced in Section 3. We will not delve into the details of the itemsets that contributed in the scores presented for each data set but we will give some basic guidelines on how the contribution matrix $\boldsymbol{\mathrm{{C}}}$ and the Nominal Outlyingness Depth can facilitate their interpretation. For instance, the contribution matrices for the outliers of each data set are illustrated in Figure 8. As can be seen, the size of the sun spot area or whether there was any activity in that region in the last twenty-four hours seem to be the most influential variables for a solar flare to be flagged as possibly moderate or severe. Similarly, the stage, metastasis (M) classification and whether the patient has received radioactive iodine therapy are indicators of thyroid cancer recurrence, while marginal adhesion and mitoses can be used to detect cases of malignant breast tumor. Finally, the rate of lymphatic diminishment and the blockage of lymphatic sinuses can hint normal finds or cases of fibrosis in the Lymphography data set. We also plot the Nominal Outlyingness Depth against the scores for the outliers in the Solar Flare data set, for which $\operatorname*{\operatorname{MAXLEN}}$ was larger than one in Figure 9. Some observations have a zero depth and score, while most outliers have at least a depth of one and one observation even has a depth of three. The largest score observed for the Solar Flare data set corresponds to a depth of nearly two units, suggesting that contributions from itemsets of both unit and greater lengths have resulted in this substantial score, making this specific observation a probable outlier.
Figure 8: Contribution matrices for UCI data sets; the contributions have been scaled to lie within the unit interval and the relative contribution is presented. Only the outlying observations are presented.
Figure 9: Nominal Outlyingness Depth (NOD) against scores of nominal outlyingness for the outliers of the Solar Flare data set.
6 Conclusion
In this paper, we propose a novel framework for quantifying outlyingness in data sets consisting of nominal variables. We define nominal outlyingness as the degree to which a sequence of nominal values and its subsets differ from a Multinomial distribution with varying event probabilities that is assumed to have generated them. Based on this definition, one may define outliers as observations consisting of highly frequent of highly infrequent sequences and supersets or subsets of sequences, respectively.
Under the framework implied by the proposed definition of nominal outlyingness, we introduce a method for quantifying outlyingness for nominal data. Our proposal includes ideas from the association rule mining literature and can be seen as a extension of the work of [24]. We present a strategy that can be used to determine values for the hyperparameters that are involved in the calculation of the score of nominal outlyingness and provide details on support-based pruning, which seeks to reduce computational complexity. Furthermore, we define the concept of nominal outlyingness depth and the contribution matrix. These can be used to assess the source of nominal outlyingness by providing the user with the average length of highly frequent/infrequent sequences included in an observation, as well as with the extent to which each nominal variable has contributed to the score of outlyingness of an observation.
Our proposed framework has been compared to two state-of-the-art algorithms for nominal outlyingness quantification, namely FPOF and FPI. The efficacy our method was illustrated via a series of simulations on 2700 synthetic data sets, where the results aligned with what was expected given the data generating process. In contrast, the two competing algorithms acted in a completely agnostic manner, inflating the scores for observations with nominal levels that were not genuinely outlying but were set by default to appear fewer times than others. We have further applied our method on four real-world data sets. Two of these data sets have not been previously used for testing outlier detection but rather for classification purposes, while the other two typically serve as benchmarking data sets for newly developed outlier detection algorithms. Our method has produced comparable and in certain instances even better results to these of FPOF and FPI, under the assumption of the minority class being the class of outliers in these data sets. Moreover, our proposed framework has selected suitable hyperparameter values without requiring user input, a feature that is not supported by the other two methods.
The proposed framework has yielded some remarkable results that can be easily interpreted (thanks to the introduction of the nominal outlyingness depth and the contribution matrix) on the real-world data. However, it does not come with no limitations. A shortcoming of the framework presented is the way the hyperparameter $\operatorname*{\operatorname{MAXLEN}}$ is computed. For instance, if we are interested in highly infrequent itemsets and we have a large number of nominal variables with highly imbalanced level proportions, the expected number of sequences containing only the most frequent levels is typically large enough to produce sparse expected contingency tables. A direct consequence of this is that $\operatorname*{\operatorname{MAXLEN}}$ will be very large, leading to potentially redundant calculations and therefore an increased computational cost. This has already been observed in some cases in our simulations, where the depth was much lower than the computed value of $\operatorname*{\operatorname{MAXLEN}}$, indicating that sequences of greater length have probably had limited to no contribution to the obtained scores of nominal outlyingness. An amendment in the way $\operatorname*{\operatorname{MAXLEN}}$ is determined would alleviate this issue and reduce the computational complexity, without changing the way the results are interpreted and the conclusions derived. Additionally, one could extend the simulation study by enforcing a dependence structure among variables which are likely to be associated in a specific manner. This way, we can get rid of the independence assumption (which is typically not valid) and potentially obtain more interesting results.
We conclude by giving some recommendations on how our framework could be used for extending the research that is currently being conducted in certain fields. The proposed framework could be used for assigning weights to observations in a data set prior to running a classification algorithm. As an example, one would perhaps want to downweigh subjects with a large score of nominal outlyingness, as these can be seen as potential outliers which may distort the classification output. This can be easily done by modifying the relevant loss function to account for observation weights. Similarly, one might wish to implement such an approach for clustering categorical data or deriving a version of the Trimmed K-Means [16] algorithm for nominal data. Finally, the proposed framework could contribute in the development of more reliable predictions in linear regression with nominal predictors. Extremities are commonly the reason behind biased parameter estimates and misleading predictions of a regression model; being able to diminish the impact of potential outliers could help mitigate these problems and develop robust regression algorithms for nominal data. These recommendations could pave the way for future research opportunities and culminate in the development of novel methodologies which will in turn provide insight on research questions of interest.
Supplementary information
The R code for the proposed method and the simulations conducted is available on GitHub by accessing the repository in: https://anonymous.4open.science/r/SONO-7CD4/.
Declarations
Conflict of interest
The authors declare that there are no conflict of interest to disclose.
Funding
The first author gratefully acknowledges funding provided by EPSRC’s Centre for Doctoral Training in Modern Statistics and Statistical Machine Learning grant EP/S023151/1.